OpenStudy (studyx3):

The length of a rectangle is 7 more than the width, and the perimeter is 38. What is the width and length?

OpenStudy (mathmale):

Choose letters to represent the length and width of the rectangle. I would choose l and w for obvious reasons. Now, write the relationship that shows how the length (l) depends upon the width (w).

OpenStudy (ben00):

Do 7 x 2 and do it minus 38?

OpenStudy (studyx3):

@mathmale would it be L+L+W+7+W+7=38?

OpenStudy (mathmale):

Your formula for the perimeter of the rectangle is a step in the right direction. You are apparently substituting W + 7 for L (which is correct). But you then have an expression for L + L + L + L, which is not correct. Think: how could you eliminate either L or W? I'd suggest you eliminate L and keep W. Start with L=W+7 and substitute this for L in the expression 2W + 2L = 38.